the area of the garden is 50 square feet, and the length of the garden is twice the width. if the fencing costs three dollars per foot, what will be the price of the fence?
Answer by ewatrrr(24785)

A =Lw = 2w(w) = 50ft^2
w^2 = 25, w = 5 (Tossing out negative Solution)
P = 2L + 2w = 2*10 + 2*5 = 30ft
Cost(fencing) = $3(30ft) = $90